Feedback and Sensitivity in Climate

Alain Lahellec June 16, 2009

Outline Make Titles Informative. Use Uppercase Letters.

Subtitles are optional.

Use itemize a lot. Use very short sentences or short phrases.

university-logo

Make Titles Informative.

You can create overlays… using the pause command:

\{ \begin{matrix}∂_{t}σ_{0}(t) = M^{♭}σ_{0}(t, 0)+|b> Υ(t) \\ g(t) =<c| σ_{0}(t, 0) , \end{matrix}  (1)

Δθ(t) =<c(t)| Φ(t, 0)´ η_{0}^{r}

+ <c| (t) ∫_{0}^{t} Φ(t, τ)|b> (τ)Δθ_{0}(τ)dτ (2)

using overlay specifications:

First item. Second item.

with the definition of the propagator as:

∂_{t}Φ(t, τ) = M(t)Φ(t, τ) ; Φ(τ, τ) = I ,  (3)

with the identity matrix I. using the general uncover command:

First item. Second item.

university-logo

You can create overlays… using the pause command:

\{ \begin{matrix}∂_{t}σ_{0}(t) = M^{♭}σ_{0}(t, 0)+|b> Υ(t) \\ g(t) =<c| σ_{0}(t, 0) , \end{matrix}  (1)

Δθ(t) =<c(t)| Φ(t, 0)´ η_{0}^{r}

+ <c| (t) ∫_{0}^{t} Φ(t, τ)|b> (τ)Δθ_{0}(τ)dτ (2)

using overlay specifications:

First item. Second item.

with the definition of the propagator as:

∂_{t}Φ(t, τ) = M(t)Φ(t, τ) ; Φ(τ, τ) = I ,  (3)

with the identity matrix I. using the general uncover command:

First item. Second item.

university-logo

You can create overlays… using the pause command:

\{ \begin{matrix}∂_{t}σ_{0}(t) = M^{♭}σ_{0}(t, 0)+|b> Υ(t) \\ g(t) =<c| σ_{0}(t, 0) , \end{matrix}  (1)

Δθ(t) =<c(t)| Φ(t, 0)´ η_{0}^{r}

+ <c| (t) ∫_{0}^{t} Φ(t, τ)|b> (τ)Δθ_{0}(τ)dτ (2)

using overlay specifications:

First item. Second item.

with the definition of the propagator as:

∂_{t}Φ(t, τ) = M(t)Φ(t, τ) ; Φ(τ, τ) = I ,  (3)

with the identity matrix I. using the general uncover command:

First item. Second item.

university-logo

You can create overlays… using the pause command:

\{ \begin{matrix}∂_{t}σ_{0}(t) = M^{♭}σ_{0}(t, 0)+|b> Υ(t) \\ g(t) =<c| σ_{0}(t, 0) , \end{matrix}  (1)

Δθ(t) =<c(t)| Φ(t, 0)´ η_{0}^{r}

+ <c| (t) ∫_{0}^{t} Φ(t, τ)|b> (τ)Δθ_{0}(τ)dτ (2)

using overlay specifications:

First item. Second item.

with the definition of the propagator as:

∂_{t}Φ(t, τ) = M(t)Φ(t, τ) ; Φ(τ, τ) = I ,  (3)

with the identity matrix I. using the general uncover command:

First item. Second item.

university-logo

You can create overlays… using the pause command:

\{ \begin{matrix}∂_{t}σ_{0}(t) = M^{♭}σ_{0}(t, 0)+|b> Υ(t) \\ g(t) =<c| σ_{0}(t, 0) , \end{matrix}  (1)

Δθ(t) =<c(t)| Φ(t, 0)´ η_{0}^{r}

+ <c| (t) ∫_{0}^{t} Φ(t, τ)|b> (τ)Δθ_{0}(τ)dτ (2)

using overlay specifications:

First item. Second item.

with the definition of the propagator as:

∂_{t}Φ(t, τ) = M(t)Φ(t, τ) ; Φ(τ, τ) = I ,  (3)

with the identity matrix I. using the general uncover command:

First item. Second item.

university-logo

You can create overlays… using the pause command:

\{ \begin{matrix}∂_{t}σ_{0}(t) = M^{♭}σ_{0}(t, 0)+|b> Υ(t) \\ g(t) =<c| σ_{0}(t, 0) , \end{matrix}  (1)

Δθ(t) =<c(t)| Φ(t, 0)´ η_{0}^{r}

+ <c| (t) ∫_{0}^{t} Φ(t, τ)|b> (τ)Δθ_{0}(τ)dτ (2)

using overlay specifications:

First item. Second item.

with the definition of the propagator as:

∂_{t}Φ(t, τ) = M(t)Φ(t, τ) ; Φ(τ, τ) = I ,  (3)

with the identity matrix I. using the general uncover command:

First item. Second item.

university-logo

ps2pdf OK.

Figure: ClimSI WV feedback model responses.

university-logo

Essai avec pgfuse.

image

Summary

The first main message of your talk in one or two lines. The second main message of your talk in one or two lines. Perhaps a third message, but not more than that.

Outlook

Something you haven’t solved. Something else you haven’t solved.

For Further Reading I

A. Author. Handbook of Everything. Some Press, 1990.

S. Someone. On this and that. Journal of This and That, 2(1):50–100, 2000.